# The Atlas Model and SDEs with Boundary Interaction

### v1

@@ 3.1 Convergence to the Generalised Mean-Field Limit
We proceed in several steps. First, we transform PDE (5.1) to a PDE on the whole space. Next, we apply a suitable time change to obtain a PDE that is driven by differentiable integrators in place of $f$ and $g$. This PDE will be within the scope of the classical superposition principle by Figalli [16, Theorem 2.6], which does most of the heavy lifting of the proof. This superposition principle provides a weak solution to a modified SDE, from which we can construct a weak solution to SDE (5.2) by reversing the transformations applied to PDE (5.1).
@@
for $\varphi\in C_{c}^{2}(\mathbb{R})$. Thus, by [16, Theorem 2.6], on a suitable filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ supporting an $\mathbb{F}$-Brownian motion $\tilde{W}$ and a real-valued $\mathcal{F}_{0}$-measurable random variable $\tilde{\xi}$ with distribution $\nu_{0}=\tilde{\mu}_{0}$, we can find an $\mathbb{F}$-adapted weak solution $\tilde{Z}=(\tilde{Z}_{t})_{t\geq 0}$ to the SDE
@@
The final step is convert the above SDE, defined on the whole space, into a reflected SDE on the half-line $0,\infty)$. For that, we define the process $X=(X_{t})_{t\geq 0}$ by $X_{t}=\lvert Z_{t}\rvert$, so that $X_{0}=\lvert Z_{0}\rvert\sim\mu_{0}$. By the Itô–Tanaka formula (cf. [23, Chapter 3, Theorem 7.1]), we have that
@@
The observations we made so far allow us to apply the superposition principle, Proposition 5.1. Thereby, $\mu_{t}=\mathcal{L}(\tilde{X}_{t})$ for $t\geq 0$, where $(\tilde{X},L)$ is the solution to the Skorokhod problem for $\xi-\alpha\ell+W$. Thus, to show that $b=\alpha\ell$ is a solution to McKean–Vlasov SDE (1.3) in the sense of Definition 1.1, it remains to prove that $\mathbb{E}[L_{t}]=\gamma t$ for $t\geq 0$. Proceeding as in [3, Lemma 1.18], we first derive the classical identity
@@
for $t\geq 0$. Let $\tilde{L}^{x}$ denote the local time of $\tilde{X}$ at $x\geq 0$, so that $\tilde{L}^{0}=2L$. Note that $\tilde{L}^{x}$ can be chosen to be continuous in $t\geq 0$ and càdlàg in $x\geq 0$. Then, by the occupation time formula (see e.g. [29, Chapter VI, Corollary 1.6]), for any $\epsilon>0$, it holds that

### v2

@@ 3.1 Convergence to the Generalised Mean-Field Limit
We proceed in several steps. First, we transform PDE (5.1) to a PDE on the whole space. Next, we apply a suitable time change to obtain a PDE that is driven by differentiable integrators in place of $f$ and $g$. This PDE will be within the scope of the classical superposition principle by Figalli [undefo, Theorem 2.6], which does most of the heavy lifting of the proof. This superposition principle provides a weak solution to a modified SDE, from which we can construct a weak solution to SDE (5.2) by reversing the transformations applied to PDE (5.1).
@@
for $\varphi\in C_{c}^{2}(\mathbb{R})$. Thus, by [undefo, Theorem 2.6], on a suitable filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ supporting an $\mathbb{F}$-Brownian motion $\tilde{W}$ and a real-valued $\mathcal{F}_{0}$-measurable random variable $\tilde{\xi}$ with distribution $\nu_{0}=\tilde{\mu}_{0}$, we can find an $\mathbb{F}$-adapted weak solution $\tilde{Z}=(\tilde{Z}_{t})_{t\geq 0}$ to the SDE
@@
The final step is convert the above SDE, defined on the whole space, into a reflected SDE on the half-line $0,\infty)$. For that, we define the process $X=(X_{t})_{t\geq 0}$ by $X_{t}=\lvert Z_{t}\rvert$, so that $X_{0}=\lvert Z_{0}\rvert\sim\mu_{0}$. By the Itô–Tanaka formula (cf. [undefv, Chapter 3, Theorem 7.1]), we have that
@@
The observations we made so far allow us to apply the superposition principle, Proposition 5.1. Thereby, $\mu_{t}=\mathcal{L}(\tilde{X}_{t})$ for $t\geq 0$, where $(\tilde{X},L)$ is the solution to the Skorokhod problem for $\xi-\alpha\ell+W$. Thus, to show that $b=\alpha\ell$ is a solution to McKean–Vlasov SDE (1.3) in the sense of Definition 1.1, it remains to prove that $\mathbb{E}[L_{t}]=\gamma t$ for $t\geq 0$. Proceeding as in [undefb, Lemma 1.18], we first derive the classical identity
@@
for $t\geq 0$. Let $\tilde{L}^{x}$ denote the local time of $\tilde{X}$ at $x\geq 0$, so that $\tilde{L}^{0}=2L$. Note that $\tilde{L}^{x}$ can be chosen to be continuous in $t\geq 0$ and càdlàg in $x\geq 0$. Then, by the occupation time formula (see e.g. [undefab, Chapter VI, Corollary 1.6]), for any $\epsilon>0$, it holds that